# LInK: Learning Joint Representations of Design and Performance Spaces through Contrastive Learning for Mechanism Synthesis

### v1

@@ 4.4 Inference Speed & Scalability
Table 3: This table shows the average inference time for each method. For SOTA methods we report the rough inference time reported by the authors (they report a range, and we report the middle of that range here), as the code for the work by Pan et al. (2023) is not publicly available. For GCP-HOLO, we run both of their variants on our hardware using an Intel i9-14900K and an RTX 4090 GPU, which provides better inference speed than the authors reported. We also run our model on the same hardware. In each case, 10 tests are run, and the average time is reported below. The percentages are reported based on the inference speed up compared to MINLP. We observe that on average, LInK is 33X faster than the fastest GCP-HOLO.
@@
LInK (Ours) 20 2000 75.18 (99.79%)
Table 3 presents the inference speed and maximum mechanism sizes each method is capable of producing within that inference time. Furthermore, we show in this table the simulation fidelity, which determines the number of points sampled for each method. We can see that LInK is not only 99.67% faster than the MINLP method, but it is also more than 95% faster than the fastest method in the state of the art capable of producing mechanisms that go beyond 4-bar and 6-bar mechanisms, GCP-HOLO. Beyond this, we see that LInK is capable of producing much more complex mechanisms with up to 20 joints and simulating with 2000 timesteps during optimization. This is in contrast with the 11 joints and 20 timesteps that the best competing method is working with while having 95% slower inference. These results demonstrate that frameworks like LInK can be very powerful for speeding up these kinds of optimization-based methods by removing the majority of the burden from the optimizer, enabling faster and more precise optimization.
@@ 4.5 Validating The Effectiveness of GHop Architecture
To validate the superiority of our GHop architecture over standard GNN architectures, we conducted a comparative performance analysis. We trained the LInK framework using baseline Graph Convolutional Network (GCN) and Graph Isomorphism Networks (GIN) models, each configured with an equivalent number of layers, hops, and hidden sizes, resulting in approximately the same parameter count: 38M for both GCN and GIN, and 31M for GHop.
@@
5 Conclusion & Future Work
In this work, we introduced LInK, a novel framework that combines contrastive learning with optimization techniques for effectively solving complex engineering design problems, with a specific focus on the path synthesis problem for planar linkage mechanisms. By leveraging a vast dataset and employing a multi-modal, transformation-invariant contrastive learning framework, LInK adeptly captures intricate physics and design representations, enabling the rapid and precise retrieval and optimization of mechanisms. This approach not only significantly expedites the search process, but also enhances precision through a hierarchical optimization algorithm. The results demonstrate the effectiveness of LInK through improvement in both speed and performance by orders of magnitude.
For future work, the adaptability and robustness of LInK open new avenues for exploring its applicability across a broader spectrum of engineering challenges beyond linkage synthesis. As demonstrated, the integration of contrastive learning-based joint design and performance space representation and optimization has the potential to facilitate advancements in areas where traditional methods falter due to the complexity or combinatorial nature of design spaces. Further research could explore the extension of LInK’s methodology to more complex mechanisms, the incorporation of dynamic simulation for moving mechanisms, and a fully end-to-end generative model.
@@ Appendix A Broader Discussion On Computational Kinematic Synthesis of Planar Linkage Mechanisms
As discussed prior most methods in kinematic synthesis fall into 3 primary categories: Numerical atlas-based approaches, Optimization-based approaches, and Deep learning-based approaches.
@@
In numerical atlas approaches first a database of mechanisms is created or gathered and simulated to obtain the paths that each mechanism traces in the database. Then these mechanisms and their corresponding paths are used as a “numerical atlas”. Given a target path, one can look up all of the mechanisms in a database to find the best matching candidate in the database McGarva (1994); Chu & Sun (2010b); Sun et al. (2015). It can be computationally expensive to compare a target to a large number of mechanisms therefore in the majority of these approaches, the database is usually limited in size or limited to a handful of mechanism types, such as a four-bar or a six-bar mechanism.
@@
A different kind of approach to solving this kind of problem is applying different optimization approaches. Some use genetic algorithms to the problem Lipson (2008); Khan et al. (2015b), While others have developed Fourier descriptor-based optimization algorithms Ullah & Kota (1997); Wu et al. (2011). At the same time, many approaches focus on specific problems without generalizability or refining existing solutionsLipson (2008); Bächer et al. (2015). Most of the aforementioned methods commonly work with existing solutions Bächer et al. (2015); Thomaszewski et al. (2014) or are limited to specific kinds of mechanisms or problems Lipson (2008); Baskar & Plecnik (2021). Most promising amongst the recent developments is the method proposed by Pan et al. (2023), which formulates the optimization problem discussed prior as a mixed integer conic-programming (MICP) and mixed-integer nonlinear programming (MINLP). However, these methods use a branch and bounds method to optimize the problem, which limits their capacity in the number of joints and the number of points in their target curves Pan et al. (2023). Furthermore, even for simple problems the time to solve any given input is significant (in the order of 5-10 hours) which makes these kinds of methods not only limited by mechanism size and target path point count but prohibitive for generating multiple candidates for the same problem or generating solutions for complex targets that require many points to describe.
@@
Given these limitations in the conventional approaches, data-driven and deep learning approaches have been proposed to accelerate and improve path synthesis results. These studies often integrate traditional methods, such as the “numerical atlas” and optimization techniques, into data-driven frameworks. For instance, Deshpande & Purwar (2019a) have refined a method that merges the numerical atlas concept with optimization strategies (Deshpande & Purwar, 2019a; b; 2020), utilizing variational autoencoders (VAEs) (Kingma & Welling, 2014) and a clustering-based search to identify suitable candidates for generating the required coupler curves. Additionally, their subsequent research applies VAEs and conditional VAEs (Sohn et al., 2015) to create mechanisms. However, the datasets employed in these studies are typically small and restricted to certain mechanism types (e.g., four-bar, six-bar, etc.), with one instance being a dataset comprising 6818 linkage mechanisms (Deshpande & Purwar, 2021). Other data-driven studies have explored mechanism generation by conditioning on specific paths (Vasiliu & Yannou, 2001), yet this approach is also mainly limited to four-bar mechanisms. In contrast to these strategies, methods are attempting to apply machine learning to emulate optimization techniques. An example of this is a study employing deep Q learning (Mnih et al., 2013). More recently Fogelson et al. (2023) have introduced a novel reinforcement learning approach that accelerates the process of optimization in comparison to MICP-based approaches mentioned before. Although these reinforcement learning-based strategies are not confined to particular mechanisms, they require retraining for each new target shape, thus limiting their generality and introducing issues similar to the aforementioned MICP-based methods.
@@ B.1 Learning Based Retrieval
When it comes to deep learning-based approaches for retrieval the majority of research has been dedicated to vision applications, specifically content-based image retrieval (CIBR). Dubey (2022) and Chen et al. (2023) have conducted comprehensive surveys on the deep learning developments for different types and applications of CIBR, which provide a more in-depth discussion on the topic for readers seeking a more detailed review of the topic. Here we will focus on discussing some of the more directly related works of research on the topic. In our work, we aim to retrieve mechanisms (represented as graphs) given target kinematics (2D paths), which makes the approach one that involves different modalities of data representation. As such here we focus on cross-modal retrieval methods, discuss some of the ways that others have approached problems of this nature, and refer readers to the aforementioned surveys for a review of other types of retrieval models.
@@
Cross-modal image retrieval refers to the task of retrieving images in a dataset based on information involving more than one modality. A prominent example of this is text and image-based retrieval, which comes up in text query-based image retrieval and image labeling based on text retrieval. One of the earliest works on the topic by Feng et al. (2014) called correspondence autoencoder (Corr-AE) trains two autoencoders simultaneously one for text and one for images and attempts to build correspondence by encouraging the embeddings of the two autoencoders to be close to each other through an L2 loss function between the text and image embeddings produced by each autoencoder. Cao et al. (2016) introduce a novel Deep Visual-Semantic Hashing (DVSH) model that introduces an end-to-end deep learning architecture called visual-semantic fusion network, which simultaneously captures spatial dependencies in images and temporal dynamics in text. This network is responsible for learning a joint embedding space that mitigates the heterogeneity between the two modalities, thereby generating compact, similarity-preserving hash codes, which can be used for text-based retrieval. Others have looked at similar approaches with slight differences in architecture and implementation such as Textual-Visual Deep Binaries (TVDB) Shen et al. (2017) and many more approaches that have been explored for this kind of cross-modal task with similar overall methodologies Wei et al. (2017); Yang et al. (2017); Wu et al. (2017) with slight nuances the detailed description of which is outside the scope of this section. Furthermore, some research has been focused on performing the same task through adversarial learning approaches. For example Li et al. (2018) introduces Self-Supervised Adversarial Hashing (SSAH). The methodology combines adversarial learning with a self-supervised semantic network to achieve this goal. Two adversarial networks are employed to learn the high-dimensional features and corresponding hash codes for different modalities. Simultaneously, a self-supervised semantic network, leveraging multi-label annotations, guides the adversarial learning process to maximize semantic relevance and feature distribution consistency between modalities. Again various works of research have been proposed along the same lines of using adversarial networks Wang et al. (2017); Zhang et al. (2018); Gu et al. (2019); Ji et al. (2019) for this kind of text-to-image or image-to-text retrieval.
@@
The common pattern across the aforementioned works and the majority of works that surround the topic is the fact that these methods all focus primarily on text and image, which means the authors often tailor their approaches and architecture to specialize in this kind of task specifically. This makes these models much better at performing cross-modal retrieval on images and text but makes the approaches very difficult to adapt to general cross-modal retrieval tasks such as the one we are performing in this work. This means that we must look at methods that are robust to cross-modal retrieval tasks which do not specialize at their core to deal specifically with image and text and this is where contrastive learning-based retrieval methods come into play.
@@ B.2 Contrastive Learning and Contrastive Learning Based Retrieval
As discussed before most of the methods built for retrieval using deep learning were built to be specialized and included mechanisms specific to text and image datasets which are not adaptable to general tasks like the one we intend to perform here. Given this and the fact that the inverse kinematic problem involves a very general and context-free nature, the kinematics simulations are the only directly available information for any given mechanism and this exists in isolation, unsupervised and self-supervised approaches are best suited for building a robust retrieval mechanism for kinematic synthesis. These requirements lend themselves rather well to contrastive learning.
@@
Contrastive learning at its core refers to unsupervised or self-supervised learning methods that aim to develop models that learn to distinguish between similar and dissimilar data points Jaiswal et al. (2021). These methods gained popularity as a result of the early seminal works in deep contrastive learning Hjelm et al. (2018); van den Oord et al. (2019); Chen et al. (2020a) that demonstrated the powerful capabilities of these approaches for both conventionally supervised deep learning problems (e.g. contrastive learning feature extraction for classification and regression) and unsupervised representation learning. Among these early approaches, the work by Chen et al. (2020a; b), known as an effective framework for contrastive learning (SimCLR) stands out as a rather generalizable and robust framework. Although SimCLR itself was built and tested on a single modality of data, primarily images, similar approaches soon emerged in other works of research that adapted similar techniques to contrastive learning on multiple modalities. Most notably the work by Radford et al. (2021), introduces the Contrastive Language-Image Pretraining (CLIP) model, which creates a crossmodal embedding space for text and images. However, unlike highly specialized text-image retrieval models the CLIP approach is rather generalizable to other cross-modal embedding spaces. This is evident in the loss function used by CLIP (Equation 2). This means that so long as trainable models can be implemented for data representations in each modality the CLIP approach applies to the problem. Given this robustness and generalizability, we incorporate the CLIP loss in our approach for unsupervised representation learning in the inverse kinematics problem.
@@
The concept of a shared cross-modal embedding space is highly analogous to the retrieval problem discussed before. This is because these models map data from different modalities into the same space based on similarity, which makes these embedding spaces ripe for retrieval problems. As it turns out, this has been explored in many works of research surrounding cross-modal retrieval. In an approach very similar to what we do here Izacard et al. (2022) uses this concept of contrastive embedding spaces for retrieval for the task of document retrieval. These kinds of approaches have also been explored for cross-modal retrieval tasks as well, such as temporal moment retrieval from videos based on text Zhang et al. (2021) or text-based molecular retrieval for molecule design Liu et al. (2023), which demonstrates the effectiveness of cross-modal embedding spaces built using contrastive learning for retrieval tasks. In our approach, we employ contrastive learning-enabled cross-modal spaces for mechanism retrieval based on kinematic objectives. The details of our implementation are discussed in later sections.
Appendix C Additional Visualizations

### v2

@@ 4.4 Inference Speed and Scalability
Table 3: This table shows the average inference time for each method. For SOTA methods we report the rough inference time reported by the authors (they report a range, and we report the middle of that range here), as the code for the work by Pan et al. (2023) is not publicly available. For GCP-HOLO, we run both of their variants on our hardware using an Intel i9-14900K and an RTX 4090 GPU, which provides better inference speed than the authors reported. We also run our model on the same hardware. In each case, 10 tests are run, and the average time is reported below. The percentages are reported based on the inference speed up compared to MINLP. We observe that, on average, LInK is 150 times faster than the fastest GCP-HOLO. At the bottom of the table, we also include the timing of manually searching the dataset of 10M samples for retrieval only using manual ordered distance calculation and using cosine search using for contrastive retrieval. The value of contrastive retrieval is clear with multiple orders of magnitude faster retrieval compared to manual search.
@@
LInK (Ours) 20 2000 15.49 (99.95%)
Manual Search of 10M Samples (GPU) 20 - 10178
Manual Search of 10M Samples (CPU) 20 - 48421
Contrastive Retrieval In 10M Samples (GPU) 20 - 0.0259
Contrastive Retrieval In 10M Samples (CPU) 20 - 6.407
Table 3 presents the inference speed and maximum mechanism sizes that each method is capable of producing within that inference time. Furthermore, we show in this table the simulation fidelity, which determines the number of points sampled for each method. We can see that LInK is not only 99.95% faster than the MINLP method, but it is also more than 99% faster than the fastest method in the state of the art – capable of producing mechanisms that go beyond 4-bar and 6-bar mechanisms, GCP-HOLO. Beyond this, we see that LInK is capable of producing much more complex mechanisms with up to 20 joints and simulating with 2000 timesteps during optimization. This is in contrast with the 11 joints and 20 timesteps that the best competing method is working with while having 99% slower inference. We also look at the inference speed of retrieval using manual search (i.e., measuring ordered distance to all curves in the dataset manually and picking the top candidates) and the contrastive retrieval approach we propose. We see clearly that manually searching the dataset would be computationally prohibitive and would require times closer to the slowest baselines of MINLP and MICP, while contrastive retrieval takes under 0.1 seconds accelerating retrieval by many orders of magnitude, while also allowing for retrieval based on partial curves which would be not be as simple as measuring ordered distance. These results demonstrate that frameworks like LInK can be very powerful for speeding up these kinds of optimization-based methods by removing the majority of the burden from the optimizer, enabling faster and more precise optimization.
@@ 4.5 Validating The Effectiveness of GHop Architecture
To validate the advantage of our GHop architecture over standard GNN architectures, we conducted a comparative performance analysis. We trained the LInK framework using baseline Graph Convolutional Network (GCN) and Graph Isomorphism Networks (GIN) models, each configured with an equivalent number of layers, hops, and hidden sizes, resulting in approximately the same parameter count: 38M for both GCN and GIN, and 31M for GHop.
@@
5 Conclusion and Future Work
In this work, we introduced LInK, a novel framework that combines contrastive learning with optimization techniques for effectively solving complex engineering design problems, with a specific focus on the path synthesis problem for planar linkage mechanisms. By leveraging a vast dataset and employing a multimodal, transformation-invariant contrastive learning framework, LInK adeptly captures intricate physics and design representations, enabling the rapid and precise retrieval and optimization of mechanisms. This approach not only significantly expedites the search process, but also enhances precision through a hierarchical optimization algorithm. The results demonstrate the effectiveness of LInK through improvement in both speed and performance by an order of magnitude.
For future work, the adaptability and robustness of LInK opens new avenues for exploring its applicability across a broader spectrum of engineering challenges, beyond linkage synthesis. Direct analogs exist in soft robotics and compliant mechanisms Sun et al. (2024), where kinematic synthesis of a similar nature is investigated in non-rigid body components. In these fields, particularly compliant mechanisms, there often exists a mixed integer representation of the design space and a curve representation of the performance space with a differentiable solver for continuous variables Sun et al. (2024). Less direct analogs can be found in complex optimization problems such as topology optimization Wang et al. (2021). For instance, in structural topology optimization for minimizing compliance (maximizing stiffness) with complex physics like buckling Ferrari & Sigmund (2019), conventional methods can take hours or days to solve. Our framework could potentially capture such physics in a cross-modal contrastive space, significantly accelerating the process of finding good local candidates. As demonstrated, the integration of contrastive learning-based joint design and performance space representation and optimization has the potential to facilitate advancements in areas where traditional methods falter due to the complexity or combinatorial nature of design spaces. Further research could explore the extension of LInK’s methodology to these fields, demonstrating its ability to handle problems that conventional optimization struggles with.
@@ Appendix A Broader Discussion On Computational Kinematic Synthesis of Planar Linkage Mechanisms
As discussed prior most methods in kinematic synthesis fall into three primary categories: numerical atlas-based approaches, otimization-based approaches, and deep learning-based approaches.
@@
In numerical atlas approaches, a database of mechanisms is first created or gathered and simulated to obtain the paths that each mechanism traces in the database. Then these mechanisms and their corresponding paths are used as a numerical atlas. Given a target path, one can look up all of the mechanisms in a database to find the best matching candidate in the database McGarva (1994); Chu & Sun (2010b); Sun et al. (2015). It can be computationally expensive to compare a target to a large number of mechanisms; therefore, in the majority of these approaches, the database is usually limited in size or limited to a handful of mechanism types, such as a four-bar or a six-bar mechanism.
@@
A different kind of approach to solving this kind of problem is applying different optimization approaches. Some apply genetic algorithms to the problem Lipson (2008); Khan et al. (2015b), While others have developed Fourier descriptor-based optimization algorithms Ullah & Kota (1997); Wu et al. (2011). At the same time, many approaches focus on specific problems without generalizability or refining existing solutionsLipson (2008); Bächer et al. (2015). Most of the aforementioned methods commonly work with existing solutions Bächer et al. (2015); Thomaszewski et al. (2014) or are limited to specific kinds of mechanisms or problems Lipson (2008); Baskar & Plecnik (2021). Most promising amongst the recent developments is the method proposed by Pan et al. (2023), which formulates the optimization problem discussed prior as a mixed-integer conic-programming (MICP) and mixed-integer nonlinear programming (MINLP). However, these methods use a branch and bounds method to optimize the problem, which limits their capacity in the number of joints and the number of points in their target curves Pan et al. (2023). Furthermore, even for simple problems, the time to solve any given input is significant (in the order of 5-10 hours) which makes these kinds of methods not only limited by mechanism size and target path point count but prohibitive for generating multiple candidates for the same problem or generating solutions for complex targets that require many points to describe.
@@
Given these limitations in the conventional approaches, data-driven and deep learning approaches have been proposed to accelerate and improve path synthesis results. These studies often integrate traditional methods, such as the numerical atlas and optimization techniques, into data-driven frameworks. For instance, Deshpande & Purwar (2019a) have refined a method that merges the numerical atlas concept with optimization strategies (Deshpande & Purwar, 2019a; b; 2020), utilizing VAEs (Kingma & Welling, 2014) and a clustering-based search to identify suitable candidates for generating the required coupler curves. Additionally, their subsequent research applies VAEs and conditional VAEs (Sohn et al., 2015) to create mechanisms. However, the datasets employed in these studies are typically small and restricted to certain mechanism types (e.g., four-bar, six-bar, etc.), with one instance being a dataset comprising 6818 linkage mechanisms (Deshpande & Purwar, 2021). Other data-driven studies have explored mechanism generation by conditioning on specific paths (Vasiliu & Yannou, 2001), yet this approach is also mainly limited to four-bar mechanisms. In contrast to these strategies, methods are attempting to apply machine learning to emulate optimization techniques. An example of this is a study employing deep Q learning (Mnih et al., 2013). More recently Fogelson et al. (2023) have introduced a novel reinforcement learning approach that accelerates the process of optimization in comparison to MICP-based approaches mentioned before. Although these reinforcement learning-based strategies are not confined to particular mechanisms, they require retraining for each new target shape, thus limiting their generality and introducing issues similar to the aforementioned MICP-based methods.
@@ B.1 Learning Based Retrieval
When it comes to deep learning-based approaches for retrieval the majority of research has been dedicated to vision applications, specifically Content Based Image Retrieval (CIBR). Dubey (2022) and Chen et al. (2023) have conducted comprehensive surveys on the deep learning developments for different types and applications of CIBR, which provide a more in-depth discussion on the topic for readers seeking a more detailed review. Here we will focus on discussing some of the more directly related works of research on the topic. In our work, we aim to retrieve mechanisms (represented as graphs) given target kinematics (2D paths), which makes the approach one that involves different modalities of data representation. As such here we focus on cross-modal retrieval methods, discuss some of the ways that others have approached problems of this nature, and refer readers to the aforementioned surveys for a review of other types of retrieval models.
@@
Cross-modal image retrieval refers to the task of retrieving images in a dataset based on information involving more than one modality. A prominent example of this is text and image-based retrieval, which comes up in text query-based image retrieval and image labeling based on text retrieval. One of the earliest works on the topic by Feng et al. (2014), called Correspondence AutoEncoder (Corr-AE) trains two autoencoders simultaneously one for text and one for images, and attempts to build correspondence by encouraging the embeddings of the two autoencoders to be close to each other through an L2 loss function between the text and image embeddings produced by each autoencoder. Cao et al. (2016) introduce a novel Deep Visual-Semantic Hashing (DVSH) model that introduces an end-to-end deep learning architecture called visual-semantic fusion network, which simultaneously captures spatial dependencies in images and temporal dynamics in text. This network is responsible for learning a joint embedding space that mitigates the heterogeneity between the two modalities, thereby generating compact, similarity-preserving hash codes, which can be used for text-based retrieval. Others have looked at similar approaches with slight differences in architecture and implementation such as Textual-Visual Deep Binaries (TVDB) Shen et al. (2017). Many more approaches that have been explored for this kind of cross-modal task with similar overall methodologies Wei et al. (2017); Yang et al. (2017); Wu et al. (2017) with slight nuances, the detailed description of which is outside the scope of this section. Furthermore, some research has been focused on performing the same task through adversarial learning approaches. For example Li et al. (2018) introduces Self-Supervised Adversarial Hashing (SSAH). The methodology combines adversarial learning with a self-supervised semantic network to achieve this goal. Two adversarial networks are employed to learn the high-dimensional features and corresponding hash codes for different modalities. Simultaneously, a self-supervised semantic network, leveraging multi-label annotations, guides the adversarial learning process to maximize semantic relevance and feature distribution consistency between modalities. Again, various works of research have been proposed along the same lines of using adversarial networks Wang et al. (2017); Zhang et al. (2018); Gu et al. (2019); Ji et al. (2019) for this kind of text-to-image or image-to-text retrieval.
@@
The common pattern across the aforementioned works and the majority of works that surround the topic is the fact that these methods all focus primarily on text and image, which means the authors often tailor their approaches and architecture to specialize in this kind of task specifically. This makes these models much better at performing cross-modal retrieval on images and text but makes the approaches very difficult to adapt to general cross-modal retrieval tasks, such as the one we are performing in this work. This means that we must look at methods that are robust to cross-modal retrieval tasks which do not specialize at their core to deal specifically with image and text, and this is where contrastive learning-based retrieval methods come into play.
@@ B.2 Contrastive Learning and Contrastive Learning Based Retrieval
As discussed before, most of the methods built for retrieval using deep learning were built to be specialized and included mechanisms specific to text and image datasets, which are not adaptable to general tasks like the one we perform here. Given this, and the fact that the inverse kinematic problem involves a very general and context-free nature, the kinematics simulations are the only directly available information for any given mechanism. This exists in isolation, unsupervised, and self-supervised approaches are best suited for building a robust retrieval mechanism for kinematic synthesis. These requirements lend themselves rather well to contrastive learning.
@@
Contrastive learning at its core refers to unsupervised or self-supervised learning methods that aim to develop models that learn to distinguish between similar and dissimilar data points Jaiswal et al. (2021). These methods gained popularity as a result of the early seminal works in deep contrastive learning Hjelm et al. (2018); van den Oord et al. (2019); Chen et al. (2020a) that demonstrated the powerful capabilities of these approaches for both conventionally supervised deep learning problems (e.g. contrastive learning feature extraction for classification and regression) and unsupervised representation learning. Among these early approaches, the work by Chen et al. (2020a; b), known as an effective framework for contrastive learning (SimCLR) stands out as a rather generalizable and robust framework. Although SimCLR itself was built and tested on a single modality of data, primarily images, similar approaches soon emerged in other works of research that adapted similar techniques to contrastive learning on multiple modalities. Most notably the work by Radford et al. (2021), introduces the Contrastive Language-Image Pretraining (CLIP) model, which creates a cross-modal embedding space for text and images. However, unlike highly specialized text-image retrieval models, the CLIP approach is rather generalizable to other cross-modal embedding spaces. This is evident in the loss function used by CLIP (Equation 2). This means that so long as trainable models can be implemented for data representations in each modality, the CLIP approach applies to the problem. Given this robustness and generalizability, we incorporate the CLIP loss in our approach for unsupervised representation learning in the inverse kinematics problem.
@@
The concept of a shared cross-modal embedding space is highly analogous to the retrieval problem discussed before. This is because these models map data from different modalities into the same space based on similarity, which makes these embedding spaces ripe for retrieval problems. As it turns out, this has been explored in many works of research surrounding cross-modal retrieval. In an approach very similar to what we do here Izacard et al. (2022) uses this concept of contrastive embedding spaces for retrieval for the task of document retrieval. These kinds of approaches have also been explored for cross-modal retrieval tasks, such as temporal moment retrieval from videos based on text Zhang et al. (2021) or text-based molecular retrieval for molecule design Liu et al. (2023), which demonstrates the effectiveness of cross-modal embedding spaces built using contrastive learning for retrieval tasks. In our approach, we employ contrastive learning-enabled cross-modal spaces for mechanism retrieval based on kinematic objectives.
Appendix C Kinematic Synthesis Problems In Planar Linkages
Figure 12: Different types of problems in inverse kinematics of planar linkage mechanisms. Note that the yellow arm is the actuator arm
In these kinds of planar linkage mechanisms, the problem of inverse kinematic synthesis can have different types of objectives, such as path synthesis, motion synthesis, and signal generation McCarthy & Soh (2010). The path synthesis problem can be described as the problem of designing linkage mechanisms that can generate a particular path that is described by a finite series of point coordinates. As such in this type of problem, the speed and timing of the motion of the mechanism are not considered as part of the objective. Motion generation on the other hand can be thought of as the generalized version of the path generation where aside from point coordinates, the orientation of an attached rigid body (such as a robot arm) and its timing and speed are also prescribed. Finally, the function generation problem refers to the problem of generating a specific series of output crank angles (or slider positions) at given angles (or positions) at the actuator, essentially transforming the signal from the actuator (angle or position) to a different signal at the output crank (or slider). See Fig. 12 for more details. LInK is created with a primary focus on the “Path Synthesis” problem, although with some tweaking the data could be adapted to be used for other types of problems such as “Function Generation” and “Motion Synthesis” (See Fig. 12).
Appendix D Additional Visualizations